i have a right triangle with a circle inside. the height is 3 and the base is 4. the long diagonal side is 5. i need to know what the radius of the circle is?
Answer by stanbon(75887)


http://mcraefamily.com/MathHelp/GeometryTriangleCenterInscribedCircle2.htm
This site describes a formula that I did not know till I went looking for it.
radius = (Area of the circle)/s, where s=(1/2)(sum of the sides)
For your problem Area = (1/2)3*4=6 sq inches.
(sum of sides)/2 = 12/2=6
So r=6/6= 1 inch
Cheers,
Stan H.